OpenStudy (anonymous):

2+2 =? Wht's 2+2!!?!

OpenStudy (uri):

2+2=4 2+2=5 2+2=fin 2+2=22

OpenStudy (anonymous):

oh that feeling when you forget the simple stuff.

OpenStudy (anonymous):

lol i feel you